OpenStudy (anonymous):

Find the derivative of: y = sq. root ( x - (x^2 - 8)^5)

OpenStudy (anonymous):

sq. root is pretty much the entire thing raised to the (1/2) power so just do that and use chain rule because its a function to the fifth power. get it?

OpenStudy (anonymous):

I've tried that and I still keep getting the wrong answer. When using the chain rule, would I keep using it until I get 2 or stop at 2x?

OpenStudy (anonymous):

Chain rule within a chain rule. \[ \frac{ d }{ dx }\sqrt{x-({x ^{2}-8}})^{5}\]

OpenStudy (anonymous):

Let u = x - ( x² - 8 )⁵ -> u' = 1 - 5 ( x² - 8 )⁴ ( 2x) Plug into: y' = u' / 2 √ u